# An SEIR network epidemic model with manual and digital contact tracing allowing delays

### v2

@@ Abstract
We consider an SEIR epidemic model on a network also allowing random contacts, where recovered individuals could either recover naturally or be diagnosed. Upon diagnosis, manual contact tracing is triggered such that each infected network contact is reported, tested and isolated with some probability and after a random delay. Additionally, digital tracing (based on a tracing app) is triggered if the diagnosed individual is an app-user, and then all of its app-using infectees are immediately notified and isolated. The early phase of the epidemic with manual and/or digital tracing is approximated by different multi-type branching processes, and three respective reproduction numbers are derived. The effectiveness of both contact tracing mechanisms is numerically quantified through the reduction of the reproduction number. This shows that app-using fraction plays an essential role in the overall effectiveness of contact tracing. The relative effectiveness of manual tracing compared to digital tracing increases if: more of the transmission occurs on the network, when the tracing delay is shortened, and when the network degree distribution is heavy-tailed. For realistic values, the combined tracing case can reduce $R_{0}$ by $20-30\%$, so other preventive measures are needed to reduce the reproduction number down to $1.2-1.4$ for contact tracing to make it successful in avoiding big outbreaks.
@@ 1 Introduction
Contact tracing stands out as one of the effective measures for controlling a pandemic. Its primary aim is to identify and interrupt the transmission chain by detecting who is spreading the infection. In this paper, we consider two contact tracing strategies: manual and digital contact tracing. The conventional (or so-called manual) contact tracing is usually performed by public health agencies, or else by self-reporting. Confirmed cases will be interviewed and asked to report potential infectious contacts, which will then be notified and tested. [ferretti_quantifying_2020] recently argued that manual contact tracing alone might not have effectively controlled the COVID-19 epidemic. Subsequently, contact tracing apps were introduced in certain countries [wymant_epidemiological_2021, app_netherlands, elmokashfi2021norway, vogt2022australia]. This digital tracing approach enhances the speed of tracing by swiftly identifying and notifying the infected contacts. Modelling studies in [kretzschmar2020impact] showed that, compared to conventional tracing approaches, digital contact tracing could have a higher effectiveness due to its inherent speed. When an app-user is diagnosed, immediate notifications are sent out to all of the app-users who have been recently in close contact with the confirmed case, recommending them to isolate and test themselves. Notably, contacts reached by manual tracing are mainly within the social circle, while digital tracing can also identify more anonymous contacts, e.g. random contacts occurring on a bus or in a store.
@@
There have been studies of modelling conventional contact tracing [ball2011threshold, ball2015stochastic, muller2016effect] and digital tracing [ferretti_quantifying_2020, kretzschmar2020impact]. This paper aims to investigate the effectiveness of combining both manual and digital tracing and the situation when only one is in place. Besides random infectious contacts focused in most papers, we also allow transmission to occur through a (social) network. Modelling contact tracing is mathematically challenging (see, e.g., an overview in [muller2021contact]). This paper focuses on the initial phase of the epidemic in a large population. The early phase of an epidemic can often be approximated by a branching process of independent individuals [ball_strong_1995]. However, the introduction of contact tracing unfortunately disrupts the independence, necessitating other modelling approaches.
@@
More specifically, we consider an SEIR epidemic spreading on a network also allowing for random contacts [keeling_networks_2005, ball2008network]; that is, individuals can be infected by their network neighbours as well as by individuals randomly chosen from the whole population. We then assume that manual tracing only happens for contacts on the network (random contacts are usually unknown), whereas digital tracing is amongst both network and random contacts, but only when both parts are app-users. Further, we introduce a tracing delay (between diagnosis and contacts being traced) for manual tracing, of which the effect has been analysed in [muller2016effect, kretzschmar2020impact]; digital contact tracing is, in contrast, triggered instantaneously.
@@ 2.1 The SEIR network epidemic model with global contacts
The contact network $G$ is characterised by a configuration model ([danon2011networks];[newman2018networks] part III, chapter 12), which features an arbitrary degree distribution $D\sim\{p_{k}\}$ ($\mathbf{P}(D=k)=p_{k}$, $k=0,1,2,...$) with finite mean $\mu$ and variance ${\sigma}^{2}$. Initially, one randomly selected index case is infectious, while the remainder of the population is susceptible. An infectious individual makes local contact with each of his/her neighbours in $G$ randomly in time through independent Poisson processes with rate $\beta_{L}$. Moreover, this infective makes global contact with other individuals (chosen randomly from the entire community) at a rate $\beta_{G}$. If the contacted individual is not susceptible, no action occurs. If susceptible, the contacted individual becomes exposed (i.e., latent) and remains latent for a random duration $T_{L}$, whose distribution is arbitrary but specified. A latent individual becomes infectious at the end of $T_{L}$. An infective remains infectious for a constant period $T_{I}\equiv\tau_{I}$, then becomes diagnosed with probability $p_{D}$; otherwise, the infective is naturally recovered. Once diagnosed or recovered naturally, individuals play no further role in the epidemic. All contact processes and the random variables describing $T_{L}$ and $D$ are assumed to be mutually independent. The epidemic stops when no latent or infectious individuals remain in the population.
@@ 2.2 Introducing manual and digital contact tracing in the model
The non-app-users can only be involved in manual contact tracing, described as follows. Once diagnosed, infected individuals are interviewed and report each of their infectee neighbours independently with probability $p_{M}$. Each reported neighbour is then tested after an independent delay time having distribution $T_{D}$ having an arbitrary but specified distribution. If testing positive the individual immediately isolate and stop spreading the infection.
App-users, on the other hand, can perform both digital and manual contact tracing. If an app-user is diagnosed, all the app-using infectees are immediately notified, tested, and isolated if infectious (stopping further transmission). Meanwhile, every non-app-using infectee neighbour is reported and traced with probability $p_{M}$ independently with independent delays $T_{D}$, just like for non-app-users. For mathematical convenience, we assume that neither manual nor digital contact tracing are iterated. That is, the traced and tested individuals of a diagnosed individual will not be further contact traced. All contact processes, reporting processes, and the random variables describing $T_{D}$ and $T_{L}$ are assumed to be mutually independent. Table 2.1 lists all the model parameters.
@@ 3.1.1 Basic reproduction number
Let $R_{0}$ denote the largest eigenvalue of $M$, then
@@ 3.2 Approximation of the early epidemic with manual and digital contact tracing
There are two types of individuals: app-users and non-app-users. While infectious, individuals have local infectious contacts with each susceptible neighbour at rate $\beta_{L}$ (and each neighbour is an app-user with probability $\pi_{A}$); have global infectious contacts with app-users at rate $\beta_{G}\pi$ and with non-app-users at rate $\beta_{G}(1-\pi)$, respectively. An infected individual’s infectious period starts after a random latent period $T_{L}$ and ends after an additional deterministic duration $\tau_{I}$. At the end of the infectious period, an individual is diagnosed with probability $p_{D}$, or else the individual recovers naturally without being diagnosed. A diagnosed non-app-user reports each of its network infectees independently with probability $p_{M}$, and each reported individual is traced after independent delays $T_{D}$. If an app-user is diagnosed, all of its app-using infectees are digitally traced immediately, and each of its local non-app-using infectees is reported with probability $p_{M}$ and manually traced after i.i.d. delays $T_{D}$. Traced individuals are tested and immediately isolate themselves if latent or infectious. Such traced individuals are however not contact traced further.
@@
In total, we would thus have $162^{4}=$ 16 types of individuals depending on the four binary categories. Among these $16$ types, $7$ are impossible: non-app-users having digital CT links (infected by local/global contacts with/without manual CT links), non-app-users that are infected by global contacts with no digital but manual CT links, app-users that are infected by global contacts with manual CT link and with/without digital CT links. This is because digital contact tracing can only occur between two app-users, and manual contact tracing happens only among local transmissions. Further, we can merge two types: one is the app-user infected on the network with both digital CT link and manual CT link; the other is the same, but without the manual CT link. The reason is that if both types of contact tracing events occur, digital contact tracing always happens first, so manual contact tracing can be neglected. Hence, there are $16-7-1=8$ types of individuals (see an illustration in Figure 3.1).
Figure 3.1: Example of an infection tree containing the eight types of individuals in Table 3.1. The circle nodes are app-users, and the squares are non-app-users. Nodes in black stand for ”diagnosed”; in white filled with diagonal lines, for ”naturally recovered”; and in grey, for ”traced”. The arrows with ”L” are local (network) infections, while those with ”G” are global infections.
Based on our assumptions, diagnosed individuals are infectious for a fixed period $\tau_{I}$, and contact tracing can only be triggered by diagnosed individuals. Following a similar argument in [ball2015stochastic], as a consequence of the deterministic infectious period $\tau_{I}$ and constant infection rates, the infection times of the infectees are independently uniformly distributed over the infectious period. Consider a diagnosed individual with $k$ infectees having CT links (either manual or digital). Label the $k$ infectees randomly and let $U_{i}$ be the time between infectee $i$ was infected until the infector is diagnosed, then $U_{1},...,U_{k}$ are hence independently and identically distributed as $U\sim U(0,\tau_{I})$. See Figure 3.2 for an illustration.
@@ 3.2.1 Calculation of the mean-offspring matrix $M^{(MD)}$
Next, we focus on individuals with manual CT links (type 2, 5). First, we let $i_{0}$ be of type 2, a non-app-user who has a manual CT link. The duration from $i_{0}$’s infection to the diagnosis of the infector follows a uniform distribution $U\sim U(0,\tau)$. If $U+T_{D}\leq T_{L}$, $i_{0}$ is latent at the time of being traced and thus has no offspring. With probability
@@
For a traced $i_{0}$ , the effective infectious period is $\max\{0,U-T_{L}+T_{D}\},$ during which the average number of global offspring is given by $\beta_{G}\mathbf{E}[(U+T_{D}-T_{L})1_{0<U+T_{D}-T_{L}<\tau_{I}}]$, where all the offspring lack CT links. The resulting expressions are $m^{(MD)}_{23}=\beta_{G}\mathbf{E}[(U+T_{D}-T_{L})1_{0<U+T_{D}-T_{L}<\tau_{I}}](1-\pi_{A})+m^{(MD)}_{13}p^{(M)}_{R}$, and $m^{(MD)}_{27}=\beta_{G}\mathbf{E}[(U+T_{D}-T_{L})1_{0<U+T_{D}-T_{L}<\tau_{I}}]\pi_{A}+m^{(MD)}_{17}p^{(M)}_{R}$, where
@@ 3.3 Approximation of the early epidemic with manual contact tracing only
Another way is that still assuming a large $n$, we could approximate the early phase of the epidemic with manual tracing by a simpler three-type branching process, denoted by $E_{M}$, with three types described in Table 3.2. Instead of considering four binary categories in Section 2.2, since digital tracing is not implemented (no app and digital CT links), we have only two binary categories: infected through the network or by global contacts, with manual CT link or not. Due to our assumption about manual tracing, those infected globally will never be reported (so no CT link). As previously discussed in Section 3.2, the infection times of infectees with manual CT links are distributed uniformly and independently on $(0,\tau_{I})$.
@@ 3.4 Approximation of the early epidemic with digital contact tracing only
where *** are some constants. Let $I_{k}$ be a $k$-by-$k$ identity matrix, the eigenvalues of $M^{(MD)}(p_{M}=0)$ solves $det(M^{(MD)}(p_{M}=0)-\lambda I_{8})=0$. By reordering the second and fifth rows to become the last two rows and similarly moving the second and fifth columns to the end, we transform the matrix $M^{(MD)}(p_{M}=0)-\lambda I_{8}$ into a block-lower-triangular form. Moreover, interchanging two rows leaves the value of the determinant unchanged, while swapping two columns only reverses the sign of the determinant. Consequently, we obtain
@@ 4.1 Parameter setting
In our numerical illustration, we focus on varying the parameters associated with contact tracing while keeping the other parameters fixed as follows. We assume the mean degree of the (social) network to equal $\mu=5$, loosely motivated by the average household size in the EU being 2.3, and each person is assumed to meet on average 3 additional people regularly [thurner2020network]. Further, we assume the deterministic infectious period to have duration $\tau_{I}=5$ (days) [kissler2020projecting], and we also assume that the latency period is deterministic with duration $T_{L}\equiv 4$ (days) [di2020impact, kissler2020projecting] having COVID-19 in mind. We assume the basic reproduction number equals $R_{0}=3$, but we also consider $R_{0}=1.5$, which could reflect a situation with physical distancing (contact rates are reduced accordingly).

### v4

@@ Abstract
We consider an SEIR epidemic model on a network also allowing random contacts, where recovered individuals could either recover naturally or be diagnosed. Upon diagnosis, manual contact tracing is triggered such that each infected network contact is reported, tested and isolated with some probability and after a random delay. Additionally, digital tracing (based on a tracing app) is triggered if the diagnosed individual is an app-user, and then all of its app-using infectees are immediately notified and isolated. The early phase of the epidemic with manual and/or digital tracing is approximated by different multi-type branching processes, and three respective reproduction numbers are derived. The effectiveness of both contact tracing mechanisms is numerically quantified through the reduction of the reproduction number. This shows that app-using fraction plays an essential role in the overall effectiveness of contact tracing. The relative effectiveness of manual tracing compared to digital tracing increases if: more of the transmission occurs on the network, when the tracing delay is shortened, and when the network degree distribution is heavy-tailed. For realistic values, the combined tracing case can reduce $R_{0}$ by 20-30%, so other preventive measures are needed to reduce the reproduction number down to $1.2-1.4$ for contact tracing to make it successful in avoiding big outbreaks.
@@ 1 Introduction
Contact tracing stands out as one of the most effective measures for controlling a pandemic. Its primary aim is to identify and interrupt the transmission chain by detecting who is spreading the infection. In this paper, we consider two contact tracing strategies: manual and digital contact tracing. The conventional (or so-called manual) contact tracing is usually performed by public health agencies, or else by self-reporting. Confirmed cases will be interviewed and asked to report potential infectious contacts, which will then be notified and tested. It has been recently argued that manual contact tracing alone might not have effectively controlled the COVID-19 epidemic [ferretti_quantifying_2020]. Subsequently, contact tracing apps were introduced in certain countries [wymant_epidemiological_2021, app_netherlands, elmokashfi2021norway, vogt2022australia]. This digital tracing approach enhances the speed of tracing by swiftly identifying and notifying the infected contacts. Modelling studies have showed that, compared to conventional tracing approaches, digital contact tracing could have a higher effectiveness due to its inherent speed [kretzschmar2020impact] . When an app-user is diagnosed, immediate notifications are sent out to all of the app-users who have been recently in close contact with the confirmed case, recommending them to isolate and test themselves. Notably, contacts reached by manual tracing are mainly within the social circle, while digital tracing can also identify more anonymous contacts, e.g. random contacts occurring on a bus or in a store.
@@
There have been studies of modelling conventional contact tracing [ball2011threshold, ball2015stochastic, muller2016effect] and digital tracing [ferretti_quantifying_2020, kretzschmar2020impact]. This paper aims to investigate the effectiveness of combining both manual and digital tracing and the situation when only one is in place. In addition to the random infectious contacts emphasized in most studies, our model also considers transmission through a social network. Modelling contact tracing is mathematically challenging (see, e.g., an overview in [muller2021contact]). This paper focuses on the initial phase of the epidemic in a large population. The early phase of an epidemic can often be approximated by a branching process of independent individuals [ball_strong_1995]. However, the introduction of contact tracing unfortunately disrupts the independence, necessitating other modelling approaches.
@@
More specifically, we consider an SEIR epidemic spreading on a network also allowing for random contacts [keeling_networks_2005, ball2008network]; that is, individuals can be infected by their network neighbours as well as by individuals randomly chosen from the whole population. We then assume that manual tracing only happens for contacts on the network (random contacts are usually unknown), whereas digital tracing is amongst both network and random contacts, but only when both parts are app-users. Further, we introduce a tracing delay (between diagnosis and contacts being traced) for manual tracing, of which the effect has been analysed in [muller2016effect, kretzschmar2020impact]; digital contact tracing is, in contrast, triggered instantaneously. In reality, digital tracing may also encounter some tracing delays, but these are generally negligible compared to the manual tracing delays [wilmink2020real]. Intuitively, longer tracing delays reduce the effectiveness of contact tracing. This is because infected individuals have more time to potentially transmit the disease before being identified and isolated.
@@ 2.1 The SEIR network epidemic model with global contacts
The contact network $G$ is characterised by a configuration model ([danon2011networks];[newman2018networks] part III, chapter 12), which features an arbitrary degree distribution $D\sim\{p_{k}\}$ ($\mathbf{P}(D=k)=p_{k}$, $k=0,1,2,...$) with finite mean $\mu$ and variance ${\sigma}^{2}$. Initially, one randomly selected index case is infectious, while the remainder of the population is susceptible. An infectious individual makes local contact with each of his/her neighbours in $G$ randomly in time through independent Poisson processes with rate $\beta_{L}$. Moreover, this infective makes global contact with other individuals (chosen randomly from the entire community) at a rate $\beta_{G}$. If the contacted individual is not susceptible, no action occurs. If susceptible, the contacted individual becomes exposed (i.e., latent) and remains latent for a random duration $T_{L}$, whose distribution is arbitrary but specified. A latent individual becomes infectious at the end of $T_{L}$. An infective remains infectious for a constant period $T_{I}\equiv\tau_{I}$, then becomes diagnosed with probability $p_{D}$; otherwise, the infective naturally recovers. Once diagnosed or recovered naturally, individuals play no further role in the epidemic. All contact processes and the random variables describing $T_{L}$ and $D$ are assumed to be mutually independent. The epidemic stops when no latent or infectious individuals remain in the population.
@@ 2.2 Introducing manual and digital contact tracing in the model
The non-app-users can only be involved in manual contact tracing, described as follows: Once diagnosed, infected individuals are interviewed and report each of their infectee neighbours independently with probability $p_{M}$. Each reported neighbour is then tested after an independent delay time having distribution $T_{D}$ having an arbitrary but specified distribution. If testing positive the individual immediately isolate and stop spreading the infection.
App-users, on the other hand, are subject to both digital and manual contact tracing. If an app-user is diagnosed, all the app-using infectees are immediately notified, tested, and isolated if infectious (stopping further transmission). Meanwhile, every non-app-using infectee neighbour is reported and traced with probability $p_{M}$ independently with independent delays $T_{D}$, just like for non-app-users. For mathematical convenience, we assume that neither manual nor digital contact tracing are iterated. That is, the traced and tested individuals of a diagnosed individual will not be further contact traced. All contact processes, reporting processes, and the random variables describing $T_{D}$ and $T_{L}$ are assumed to be mutually independent. Table 2.1 lists all the model parameters.
@@ 3.1.1 Basic reproduction number
Let $R_{0}$ denote the largest eigenvalue of the mean-offspring matrix $M$ with elements $m_{ij}$ derived above, then
@@ 3.2 Approximation of the early epidemic with manual and digital contact tracing
There are two types of individuals: app-users and non-app-users. While infectious, individuals have local infectious contacts with each susceptible neighbour at rate $\beta_{L}$ (and each neighbour is an app-user with probability $\pi_{A}$); have global infectious contacts with app-users at rate $\beta_{G}\pi_{A}$ and with non-app-users at rate $\beta_{G}(1-\pi_{A})$, respectively. An infected individual’s infectious period starts after a random latent period $T_{L}$ and ends after an additional deterministic duration $\tau_{I}$. At the end of the infectious period, an individual is diagnosed with probability $p_{D}$, or else the individual recovers naturally without being diagnosed. A diagnosed non-app-user reports each of its network infectees independently with probability $p_{M}$, and each reported individual is traced after independent delays $T_{D}$. If an app-user is diagnosed, all of its app-using infectees are digitally traced immediately, and each of its local non-app-using infectees is reported with probability $p_{M}$ and manually traced after i.i.d. delays $T_{D}$. Traced individuals are tested and immediately isolate themselves if latent or infectious. Such traced individuals are however not contact traced further.
@@
In total, we would thus have $16$ types of individuals depending on the four binary categories. Among these $16$ types, $7$ are impossible: non-app-users having digital CT links (infected by local/global contacts with/without manual CT links), non-app-users that are infected by global contacts with no digital but manual CT links, app-users that are infected by global contacts with manual CT link and with/without digital CT links. This is because digital contact tracing can only occur between two app-users, and manual contact tracing happens only among local transmissions. Further, we can merge two types: one is the app-user infected on the network with both digital CT link and manual CT link; the other is the same, but without the manual CT link. The reason is that if both types of contact tracing events occur, digital contact tracing always happens first, so manual contact tracing can be neglected. Hence, there are $16-7-1=8$ types of individuals (see an illustration in Figure 3.1).
Figure 3.1: Example of an infection tree containing the eight types of individuals in Table 3.1. The circle nodes are app-users, and the squares are non-app-users. Nodes in black stand for “diagnosed”; in white filled with diagonal lines, for “naturally recovered”; and in grey, for “traced”. The arrows with “L” are local (network) infections, while those with “G” are global infections.
Based on our assumptions, diagnosed individuals are infectious for a fixed period $\tau_{I}$, and contact tracing can only be triggered by diagnosed individuals. Following a similar argument in [ball2015stochastic], as a consequence of the deterministic infectious period $\tau_{I}$ and constant infection rates, the infection times of the infectees are independently uniformly distributed over the infectious period. Consider a diagnosed individual with $k$ infectees having CT links (either manual or digital). Label the $k$ infectees randomly and let $U_{i}$ be the time between infectee $i$ was infected until the infector is diagnosed, then $U_{1},...,U_{k}$ are hence independently and identically distributed as $U\sim U(0,\tau_{I})$. See Figure 3.2 for an illustration. If the infectious period $T_{I}$ is random, $U_{i}$ will no longer be uniformly distributed and, more importantly, they will become dependent which disrupts the branching process approximation.
@@ 3.2.1 Calculation of the mean-offspring matrix $M^{(MD)}$
Next, we focus on individuals with manual CT links (type 2, 5). First, we let $i_{0}$ be of type 2, a non-app-user who has a manual CT link. The duration from $i_{0}$’s infection to the diagnosis of the infector follows a uniform distribution $U\sim U(0,\tau_{I})$. If $U+T_{D}\leq T_{L}$, $i_{0}$ is latent at the time of being traced and thus has no offspring. With probability
@@
For a traced $i_{0}$, the effective infectious period is $\max\{0,U-T_{L}+T_{D}\},$ during which the average number of global offspring is given by $\beta_{G}\mathbf{E}[(U+T_{D}-T_{L})1_{\{0<U+T_{D}-T_{L}<\tau_{I}\}}]$, where all the offspring lack CT links. The resulting expressions are $m^{(MD)}_{23}=\beta_{G}\mathbf{E}[(U+T_{D}-T_{L})1_{\{0<U+T_{D}-T_{L}<\tau_{I}\}}](1-\pi_{A})+m^{(MD)}_{13}p^{(M)}_{R}$, and $m^{(MD)}_{27}=\beta_{G}\mathbf{E}[(U+T_{D}-T_{L})1_{\{0<U+T_{D}-T_{L}<\tau_{I}\}}]\pi_{A}+m^{(MD)}_{17}p^{(M)}_{R}$, where
@@ 3.3 Approximation of the early epidemic with manual contact tracing only
Another way, still assuming a large $n$, we could approximate the early phase of the epidemic with manual tracing by a simpler three-type branching process, denoted by $E_{M}$, with three types described in Table 3.2. Instead of considering four binary categories in Section 2.2, since digital tracing is not implemented (no app and digital CT links), we have only two binary categories: infected through the network or by global contacts, with manual CT link or not. Due to our assumption about manual tracing, those infected globally will never be reported (so no CT link). As previously discussed in Section 3.2, the infection times of infectees with manual CT links are distributed uniformly and independently on $(0,\tau_{I})$.
@@ 3.4 Approximation of the early epidemic with digital contact tracing only
where * are some constants. Let $I_{k}$ be a $k$-by-$k$ identity matrix, the eigenvalues of $M^{(MD)}(p_{M}=0)$ solves $det(M^{(MD)}(p_{M}=0)-\lambda I_{8})=0$. By reordering the second and fifth rows to become the last two rows and similarly moving the second and fifth columns to the end, we transform the matrix $M^{(MD)}(p_{M}=0)-\lambda I_{8}$ into a block-lower-triangular form. Moreover, interchanging two rows leaves the value of the determinant unchanged, while swapping two columns only reverses the sign of the determinant. Consequently, we obtain
@@ 4.1 Parameter setting
In our numerical illustration, we focus on varying the parameters associated with contact tracing while keeping the other parameters fixed as follows. We assume the mean degree of the (social) network to equal $\mu=5$, loosely motivated by the average household size in the EU being 2.3 (it is important to note, however, that we do not model households explicitly), and each person is assumed to meet on average 3 additional people regularly [thurner2020network]. Further, we assume the deterministic infectious period to have duration $\tau_{I}=5$ (days) [kissler2020projecting], and we also assume that the latency period is deterministic with duration $T_{L}\equiv 4$ (days) [di2020impact, kissler2020projecting] having COVID-19 in mind. We assume the basic reproduction number equals $R_{0}=3$, but we also consider $R_{0}=1.5$, which could reflect a situation with physical distancing (contact rates are reduced accordingly).